OpenStudy (anonymous):

. Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

OpenStudy (anonymous):

to find the next numbers in the sequence you must divide by three

OpenStudy (anonymous):

What do you mean? Divide each number by 3?

OpenStudy (anonymous):

for example, if you divide 648 by 3 you get 216, divide 216 by 3 and you get 72 and so on

OpenStudy (anonymous):

24/3 = 8/3 = 2.7/3 - 0.9 , are those correct

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

So then the rule is, number divided by 3?